A pipe that is 325 ft. long is cut into three pieces. The first piece is one foot shorter then twice the second, and the third piece is four times as long as the other two pieces combined. How long are each of the pieces.

thanks!!! (i'm going to hit my limit today) lol

Answer by xcentaur(357)

let the length of the second piece be x
then length of the first is (2x-1)
then the length of the third is (2x-1+x)*4 = 4(3x-1)


now all the pieces add up to 325,so we get
x+(2x-1)+4(3x-1)=325
3x-1+4(3x-1)=325
1(3x-1)+4(3x-1)=325
5(3x-1)=325
(3x-1)=65
3x=65+1=66
x=66/3=22


hence the length of the pcs are: 43,22,260 feet resp


hope this helps,
Prabhat